OpenStudy (anonymous):

What is the graph of one cycle of a sine curve with the given characteristics? Using the form y = a sin b theta, what is an equation for the sine curve Ampltidue = 4, period 1/2pi and a < 0

OpenStudy (anonymous):

If your amplitude is four, but a is less than zero, we have to use -4 for a. If our period is usually 2pi, and in this case it is pi/2, everything has to go four times as fast, so b is four. That will give us a function \[f(\theta)=-4\sin (4 \theta)\]